Nicole:

Write the equation of a line in point-slope form that is parallel to the line y = 3x +6 and passes through the point (-5, 8)

Nicole:

@AngeI

Nicole:

@Gdeinward

Nicole:

I got: y-8=3(x-(-5) is that right?

AngeI:

You would make 6 into c for a constant You would have to replace x and y with (-5,8) 8=3(-5)+c 8=-15+c C=23 Y=3x+23

Nicole:

but its asking point slope form not slope intercept

AngeI:

OH My bad Sorry e.e 1 sec cx

Nicole:

Its okay lol

Nicole:

Did you get anything? @AngeI

AngeI:

Sorry my teacher was explaining something What you said looks right

AngeI:

Am currently in class x.x

Nicole:

Its okay https://prnt.sc/pmqfx4

Nicole:

@Gdeinward @AngeI ?

AngeI:

Please limit each post to one question each It just makes it less of a mess for everyone c: